the leader of association of jeepney drivers claims that the average daily take home pay of all jeepney drivers in pasay city is php 400. a random sample of 100 jeepney drivers in pasay city was interviewed and the average daily take home pay of these drivers is found to be php 425.use a 0.05 significance level to find out if the average daily take home pay of all jeepney drivers in pasay city is different from php 400. assume that the population variance is php8464.


1
Expert's answer
2021-06-15T14:09:14-0400

H0: "\mu=400"

Ha: "\mu\ne400"


"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{425-400}{\sqrt{8464}/\sqrt{100}}=2.72"


p-value:

"2P(z<2.72)=2\cdot0.00326=0.00652"

Since p-value < 0.05, we reject the null hypothesis. So, the average daily take home pay of all jeepney drivers in pasay city is different from php 400.